# Topological protection breakdown: a route to frustrated ferroelectricity

### v2

@@ II.3 The effect of a bias electric field
Response to a bias electric field is analyzed theoretically by adding the term $\mathcal{H}_{e}=-\sum_{i}\mathbf{h}\cdot\mathbf{p}_{i}$ in Eq. (1). As reported in Fig.2(e), the temperature versus electric field phase diagram presents a rich phenomenology. At low temperatures, the vortex phase is separated from the ferroelectric phase by a first-order phase transition, while the critical temperature $T_{p}$ gives rise to the paraelectric phase (dashed lines). Far below this temperature, a different metastable phase emerges in contrast to the previous case. As before, the order parameters ($P$ and $v$) characterize the phase transitions as a function of the external field, as shown in Figs.2(f) and (g).
@@
To establish whether the TBM actually supports a percolative phase transition, we analyzed each unit belonging to the microscopic scale, $\Lambda_{1}$, on a coarser scale: each block-dipole now represents a new node that either allows light to pass ($\ell=1$) or does not ($\ell=0$). This embodies the idea that for the KTN:Li, only the SC vortex phase allows complete light transmission. The giant cluster $P_{\infty}$ characterizes the phase transition, together with the susceptibility $\chi(h)=\frac{\sum_{S}S^{2}P(S,h)}{\sum_{S}SP(S,h)}$. Here, the sum runs over all possible cluster sizes $S$ for a given field $h$ in the system, and $P(S,h)$ the cluster size probability distribution, discarding $P_{\infty}$, if it exists.
@@
Fig. 4(a) shows the percolation phase transition at low temperatures, which confirms that an abrupt phase transition occurs at a specific critical field, $E_{c}$. Fig. 4(b) illustrates the behavior in the metastable phase, shedding light on two key properties: (i) the phase transition in the metastable region is found to be smoother, and (ii) susceptibility diverges throughout the entire region. Fig. 4(c) shows different states when the external field increases. At low temperatures, labyrinthine-like structures form where ferroelectric domains start to aggregate, giving rise to a giant cluster that spans the entire system in some critical field $E_{C}$. This leads to the metastable phase, where both the vortex and ferroelectric mesoscopic domains coexist in an interlocked way forming fractal patterns (see Fig. 4(c)).
@@
Figure 4: Percolation phase transition. (a-b) $P_{\infty}$ as a function of the applied field, $E$, for different system sizes (in the new square lattice of side $L$ resulting from the coarse-graining process, see legend) for: (a) $T=0.4$ Inset: $\chi$ versus $h$. Note the divergence at a specific point as the system size increases. (b) $T=0.75$. Inset: $\chi$ versus $h$. Note the divergence of the entire metastable region as the system size increases. (c) Configurations of simulated light transmission images. From up to down, images show a Vortex state ($T=0.4,~{}E=0.2$), the bifurcation line to the metastable phase ($T=0.5,~{}E=0.29$), and the metastable phase ($T=0.5,~{}E=0.3$). For the sake of comparison, the inset shows a zoomed region of experimental light transmission at $T_{C}$-2.5K and $E=2.65$ kV/cm.
@@
Macroscopic response is analyzed in Fig. 5, where the behavior of $P_{\infty}$ versus applied field is reported. The orange loop refers to temperature where the metastable phase is absent, leading to a strongly hysteretic reponse function typical of a first-order transition. In turn, a slim loop (violet loop) emerges in the metastable phase, whith its signature negligible hysteresis and marked nonlinearity. Characteristic discrete jumps in polarization (see the zoomed inset in Fig. 5) are caused by the depinning of finite domains, the so-called Barkhausen jumps in hysteretic loops [72, 26].
@@
Figure 5: Slim loop hysteresis. $P_{\infty}$ as a function of the applied field, $h$, for different temperatures, $T=0.4$ (orange line) and $T=0.6$ (violet line) for a lattice with $L=64$ coarse-grained units. The wide hysteretic behavior at low temperatures, characteristic of the first-order phase transition, becomes a slim loop hysteresis cycle in the intermediate phase. The break of domain walls between the pinning sites originates the jumps in the hysteretic cycle shown in the zoomed insets, giving place to the well-known Barkhausen noise.
@@ III Discussion
Complex behavior in functional materials can lead to exotic and useful phenomena, such as non-ergodicity and giant response. The driving mechanism is commonly thought to be an underlying disorder that can, for example, influence a conventional phase transition by introducing competing interaction terms between multiple components and lead to frustration [73]. In ferroelectrics, complex behavior is typically attributed to alterations in what would normally be considered a conventional paraelectric-ferroelectric structural phase transition caused by built-in compositional disorder. Disordered ferroelectrics can be modelled as a system dominated by mesoscopic polar nanoregions (PNRs), regions originally thought to have randomly distributed dipole moments driven by exchange interaction [74, 75, 26], that have more recently been shown to manifest a hierarchical structure with no sudden boundaries [76]. Complex behavior could then be explained as an analog of Griffiths phases theoretically predicted long ago (but not yet experimentally found) for dilute ferromagnetic systems [77]. Unfortunately, this approach has been demonstrated to be insufficient at temperatures lower than the Curie temperature ($T_{C}$) because it fails to consider PNR-interaction [27]. The existence of a glassy state has also been proposed as an alternative explanation [78, 29]. Glassy approaches, in turn, encounter difficulties in describing some of the basic observed features, such as the occurrence of Barkhausen jumps during the poling process, jumps that are incompatible with the continuous and monotonous glassy reorientation of dipoles on sub-micrometer length scales [26]. A complete understanding of emergent relaxor behavior in disordered ferroelectrics remains an open problem [79].
@@
Our study suggests an altogether different picture. Here, built-in disorder is relegated to generating the SC-tiling and plays no direct role in the TBM. Yet, remarkably, the approach is able to reproduce signature and hereto little understood complex and non-ergodic behavior of ferroelectrics. For one, the percolative transition typical of relaxor-like perovskites [70, 24, 80], is predicted by the TBM at low temperatures (below $T_{c}$) as an abrupt phase transition that splits into two critical electric fields with an intermediate behavior that resembles that of the metastable phase, as discussed in Figs.3 and 4). Moreover, as discussed in Fig.5, the TBM is able to describe the anomalous hysteretic ”slim-loop” behavior observed in relaxors behavior, along with the characteristic Barkhausen jumps [26]. Results indicate that complexity-driven phenomenology originates from the competition between the two intrinsically different scales ($\Lambda_{1}$ and $\Lambda_{2}$) that arise not out of disorder but geometrically, from the interaction terms of Eq.(1), as discussed in Fig.1, leading to local frustration mechanisms that persist even in the thermodynamic limit.
@@
Finally, we note that the TBM addresses the issue of describing ferroelectric behavior dominated by compositional disorder that leads to interacting mesoscopic domains relevant to optical propagation. In this regime, structural analysis and molecular dynamics simulations can prove challenging. As a qualitative model, the TBM isolates the key mechanisms behind the emergence of complex domain phases in functional materials. In line with efforts to formulate minimal yet effective theories that capture essential physics with maximal simplicity [81], the TBM reveals that it is the competition between effective dipole-dipole-like interactions and the intrinsic two-scale lattice geometry that drives the formation of such intricate domain structures.
@@ IV Conclusions
Concluding, we have introduced and experimentally validated a new explanatory framework for complex behavior observed in disordered ferroelectric crystals. Emergent phenomena are found to be a consequence of a topological symmetry-breaking associated to the competition between two basic topological defects that emerge from the spontaneous breaking of discrete inversion symmetry and volume-charge screening: domain walls parallel to the principal crystal axes, across which the polarization flips by 180 degrees, and slanted at 45 degrees, across which the polarization abruptly rotates by 90 degrees [82, 25, 36]. Frustration is then triggered by the interplay between the two corresponding spatial scales, each with its specific dipole-dipole interaction, a Heisenberg and a DM-like interaction. The mechanism leads to vortex and metastable phases, depending on the ratio of the two underlying interaction components. The application of the real-space renormalization group methods to detect domain patterns reveals the existence of a frustrated percolation phase transition that naturally includes depinning effects, i.e., domain wall reorientation that results in the well-known Barkhausen jumps and slim-loop hysteretic phenomena, when topologically protected states are broken. Our expansion of classical frustration mechanisms, beyond spin-glass and Griffiths effects, opens a new route to explore the intriguing interplay between structure and dynamics and how this can give rise to previously unknown phases, potentially rising to a new level the design and development of innovative miniaturized noise-resistant high-performance information and energy storage devices.
@@ Acknowledgments
L. F. and E. DelRe acknowledge support from PNRR MUR Project No. PE0000023-NQSTI, PRIN 2022 MUR Project No. 20223T577Z, Sapienza-Ricerca di Ateneo Projects. P.V. acknowledges the Spanish Ministry of Research and Innovation and Agencia Estatal de Investigación (AEI), MICIN/AEI/10.13039/501100011033, for financial support through Project PID2023-149174NB-I00, funded also by European Regional Development Funds, as well as Ref. PID2020-113681GB-I00. We thank A. Gabrielli, for extremely valuable discussions and suggestions on earlier manuscript versions.
@@ Appendix B Exploring the Metastable Phase
For completeness, we present here numerical results for the temporal evolution of the polarization order parameter, $P$.
@@
Our results reveal the existence of such a metastable regime placed between the vortex and the ferroelectric phase. This is characterized, as shown in Fig. 6, by broad quasi-periodic temporal oscillations of $P$. Anomalously large sampling times would be required to extract good statistics for the actual mean values and variances. This collective behavior is a straightforward manifestation of the partial ordering of the system, being reminiscent of other types of non-ergodic behavior in structured systems [33, 83] but with a completely different origin.
@@ Appendix C Experimental Setup
Laser light from a doubled 30 mW Nd:YAG laser (wavelength $\lambda=2\pi/k_{0}=532$ nm) is made to propagate along the z-axis through the zero-cut KTN:Li crystal sandwiched in between two crossed polarizers. The sample is biased by a time-constant electric field E along the x-axis. The crystal temperature T is set by a current-controlled Peltier junction in contact with one of the y facets. Light from the crystal output facet is imaged using a moveable spherical lens L1 (of focal length $50$mm) onto a CMOS camera. The cross-polarizer technique implemented utilizes the birefringence of polar clusters to alter the polarization state of light, allowing optical observation of microscopic cluster dynamics at micrometer scales. The experimental setup is illustrated in Fig. 8.
@@ Appendix D Material
The sample is a zero-cut optical quality polished lithium-enriched solid solution of potassium-tantalate-niobate (KTN: Li) with an average composition $K_{0.964}$i0.036Ta0.60Nb0.40O3. Its dimensions are 2.33(x) $\times$ 1.96(y) $\times$ 2.03(z) mm. The unit cell manifests random substitutions, a compositional disorder that, as a consequence of the structural flexibility typical of perovskites, leads to locally modified polarizabilities and temperature-dependent nanoscale dipolar structures (nanodisordered ferroelectricity). The result is a modified ferroelectric behavior dominated by the so-called polar nanoregions (PNRs), characterized by dielectric dispersion and out-of-equilibrium behavior (relaxor ferroelectricity) [35]. In our present case, this disorder is itself not homogeneous, manifesting a spatially periodic micrometric oscillation along a specific crystal axis. This is because the sample is grown into a bulk through the top-seeded method, a technique that entails a slight time oscillation in the temperature of the solidifying melt that, in turn, translates into an approximatively periodic 7.5 $\mu$ m pattern along the growth (or pull) axis, which causes in ferroelectrics an equally patterned $T_{c}$. This pattern conditions the nanoscale dipolar structures that, at the room-temperature Curie point $T_{c}$= 295K, form an orderly multi-domain protected regular state, the Supercrystal (SC) [68, 71, 70, 28]. SCs are thought to originate in a regular lattice of interlocked polarization vortices, a generalization of the spontaneous polarization closed-flux vortex domain pattern dominated by a triangular tiling, compatible with available experimental results and phase-field simulations [28]. This arrangement effectively screens the polarization charge and elastic stress [68, 84]. SC phase also manifests striking optical properties, such as giant broadband refraction (GR) [66, 69] and constraint-free wavelength conversion [67].

### v4

@@ II.3 The effect of a bias electric field
Response to a bias electric field is analyzed theoretically by adding the term $\mathcal{H}_{e}=-\sum_{i}\mathbf{E}\cdot\mathbf{p}_{i}$ in Eq. (1). As reported in Fig. 2(e), the temperature versus electric field phase diagram presents a rich phenomenology. At low temperatures, the vortex phase is separated from the ferroelectric phase by a first-order phase transition, while the critical temperature $T_{p}$ gives rise to the paraelectric phase (dashed lines). Far below this temperature, a different metastable or frustrated phase emerges in contrast to the previous case. As before, the order parameters ($P$ and $v$) characterize the phase transitions as a function of the external field, as shown in Figs. 2(f) and (g).
@@
To establish whether the TBM actually supports a percolative phase transition, we analyzed each unit belonging to the microscopic scale, $\Lambda_{1}$, on a coarser scale: each block-dipole now represents a new node that either allows light to pass ($\ell=1$) or does not ($\ell=0$). This embodies the idea that for the KTN:Li, only the SC vortex phase allows complete light transmission. The giant cluster $P_{\infty}$ characterizes the phase transition, together with the susceptibility $\chi(h)=\frac{\sum_{S}S^{2}P(S,h)}{\sum_{S}SP(S,h)}$. Here, the sum runs over all possible cluster sizes $S$ for a given field $h$ in the system, being $P(S,h)$ the cluster size probability distribution, and discarding $P_{\infty}$, if it exists.
@@
Figure 4(a) shows the percolation phase transition at low temperatures, which confirms that an abrupt phase transition occurs at a specific critical field, $E_{c}$. Fig. 4(b) illustrates the behavior in the metastable phase, shedding light on two key properties: (i) the phase transition in the metastable region is found to be smoother, and (ii) susceptibility diverges throughout the entire region. Fig. 4(c) shows different states when the external field increases. At low temperatures, labyrinthine-like structures form where ferroelectric domains start to aggregate, giving rise to a giant cluster that spans the entire system in some critical field $E_{C}$. This leads to the frustrated phase, where both the vortex and ferroelectric mesoscopic domains coexist in an interlocked way, forming fractal patterns (see Fig. 4(c)).
@@
Figure 4: Percolation phase transition. (a-b) $P_{\infty}$ as a function of the applied field, $E$, for different system sizes (in the new square lattice of side $L$ resulting from the coarse-graining process, see legend) for: (a) $T=0.4$ Inset: $\chi$ versus $E$. Note the divergence at a specific point as the system size increases. (b) $T=0.75$. Inset: $\chi$ versus $E$. Note the divergence of the entire metastable region as the system size increases. (c) Configurations of simulated light transmission images. From up to down, images show a Vortex state (T=0.4,E=0.2T=0.4,~E=0.2), the bifurcation line to the metastable phase (T=0.5,E=0.29T=0.5,~E=0.29), and the metastable phase (T=0.5,E=0.3T=0.5,~E=0.3). For the sake of comparison, the inset shows a zoomed region of experimental light transmission at $T_{C}$-2.5K and $E=2.65$ kV/cm.
@@
Macroscopic response is analyzed in Fig. 5, where the behavior of $P_{\infty}$ versus applied field is reported. The orange loop refers to temperature where frustration is absent, leading to a strongly hysteretic reponse function typical of a first-order transition. In contrast, a slim loop (violet loop) emerges in the metastable phase, a signature of negligible hysteresis and marked nonlinearity. Characteristic discrete jumps in polarization (see the zoomed inset in Fig. 5) are caused by the depinning of finite domains, the so-called Barkhausen jumps in hysteretic loops [58, 7].
@@
Figure 5: Slim loop hysteresis. $P_{\infty}$ as a function of the applied field, $E$, for different temperatures, $T=0.4$ (orange line) and $T=0.6$ (violet line) for a lattice with $L=64$ coarse-grained units. The wide hysteretic behavior at low temperatures, characteristic of the first-order phase transition, becomes a slim loop hysteresis cycle in the intermediate phase. The break of domain walls on pinned sites originates the observed jumps in the hysteretic cycle, that are shown in the zoomed insets. This gives rise to the well-known Barkhausen noise.
@@ III Discussion
A knot in a string is a remarkably stable physical form of information storage since it is topologically protected, that is, it can only be erased by untangling it [59, 60, 61, 62]. Knot-like structures emerge naturally in the form of topological defects in a wide variety of materials [63], such as liquid-crystals [64], fluids [65], ferromagnets [66, 67], and ferroelectrics [13, 14]. This protection is responsible for the stability of intricate low-dimensional topological polar structures (e.g., quadrant domains [68], polar flux closures [16, 69, 70], and vortices [71]) that, in recent years, have attracted significant attention[11, 12, 72]. Analogous topologically protected patterns naturally occur in a wide variety of systems [63], such as liquid-crystals [64], fluids [65], and ferromagnets [66, 67]. Among these, topologically protected ferromagnetic skyrmions [73] have found applications in electromechanical and spintronic devices for information storage [74] and logic [75].
@@
At present, the potential of topological defects as noise-resistant memory and processing elements is largely unexplored, especially because their behavior appears intimately connected with an intriguing but little-understood frustration in the hosting material [76, 15, 77]. In some respects, frustration likens functional materials, such as ferroelectric crystals, to the large class of information-driven phenomena, such as memory effects [78], protein-folding mechanisms [79], brain dynamics [45, 80], and slow relaxation dynamics [81]. The driving mechanism for frustration is commonly thought to be an underlying disorder that can, for example, influence a conventional phase transition by introducing competing interaction terms between multiple components and lead to frustration [4]. In ferroelectrics, complex behavior is typically attributed to alterations in what would normally be considered a conventional paraelectric-ferroelectric structural phase transition caused by built-in compositional disorder. Disordered ferroelectrics can be modelled as a system dominated by mesoscopic polar nanoregions (PNRs), regions originally thought to have randomly distributed dipole moments driven by exchange interaction [82, 83, 7], that have more recently been shown to manifest a hierarchical structure with no sudden boundaries [20]. Complex behavior could then be explained as an analog of Griffiths phases theoretically predicted long ago (but not yet experimentally found) for dilute ferromagnetic systems [84]. This approach has been demonstrated to be insufficient at temperatures lower than the Curie temperature ($T_{C}$) because it fails to consider PNR-interaction [2]. The existence of a glassy state has also been proposed as an alternative explanation [85, 1]. Glassy approaches, in turn, encounter difficulties in describing some of the basic observed features, such as the occurrence of Barkhausen jumps during the poling process, jumps that are incompatible with the continuous and monotonous glassy reorientation of dipoles on sub-micrometer length scales [7]. A complete understanding of emergent relaxor behavior in disordered ferroelectrics remains an open problem [86].
Our study suggests an altogether different picture. Here, built-in disorder is relegated to generating the SC-tiling and plays no direct role in the TBM. Yet the approach is able to reproduce the signature and hereto little-understood complex and non-ergodic behavior of disordered ferroelectrics like KTN.
@@
For one, the percolative transition typical of relaxor-like perovskites [56, 77, 87], is predicted by the TBM at low temperatures (below $T_{c}$) as an abrupt phase transition that splits into two critical electric fields with an intermediate behavior that resembles that of the new frustrated phase presented here (as discussed in Figs. 3 and 4). We note that while the cubic-to-tetragonal transition model used here is a useful paradigm for many perovskite ferroelectrics, other systems like PMN and BTO may exhibit more complex or distinct symmetry-breaking behavior. Nonetheless, our model captures the general interplay between discrete inversion-symmetry breaking and charge-screening-driven topology, which remains relevant across a wide class of materials. Moreover, as discussed in Fig. 5, the TBM is able to describe the anomalous hysteretic ”slim-loop” behavior observed in relaxor behavior, along with the characteristic Barkhausen jumps [7]. Our results indicate that complexity-driven phenomenology originates from the competition between the two intrinsically different scales ($\Lambda_{1}$ and $\Lambda_{2}$) that arise not out of disorder but geometrically, from the interaction terms of Eq.(1), as discussed in Fig.1, leading to local energetic incompatibilities that persist even in the thermodynamic limit.
Hence, the interplay between the Heisenberg-like and DMI-like terms naturally introduces frustration into the system. These two interactions, effective at different spatial scales, cannot be simultaneously minimized. We want to emphasize that other types of interactions, grounded in different topological scales as supported by the SC lattice, may also lead to similar effects. As a result, the system is unable to achieve a globally ordered state, leading to the emergence of metastable regimes characterized by local competition between polarized and vortex domains. This frustration is intrinsic to the TBM and forms the basis for the rich phase behavior discussed here.
@@
Finally, we note that the TBM addresses the issue of describing ferroelectric behavior dominated by compositional disorder that leads to interacting mesoscopic domains relevant to optical propagation. In this regime, structural analysis and molecular dynamics simulations can prove challenging. Indeed, the TBM is not intended to fully replace detailed microscopic simulations, such as phase-field or ab initio approaches. Rather, it provides a minimal and analytically tractable framework to explore how competing mesoscopic interactions and geometric constraints can produce complex, non-ergodic behavior. The simplicity of the model allows insight into qualitative mechanisms that may remain hidden in fully numerical treatments. Qualitatively, the TBM isolates key mechanisms that can lead to complex domain phases in functional materials. In line with efforts to formulate minimal yet effective theories that capture essential physics with maximal simplicity [88], the TBM reveals that it is the competition between effective dipole-dipole-like interactions and the intrinsic two-scale lattice geometry that drives the formation of such intricate domain structures. While KTN:Li is used here as a case study, the TBM is generic and applicable to other polar materials with flux-closure behavior and competing dipolar interactions that manifest superlattices. Future studies could explore applications to, e.g., SBN or multiferroic materials.
@@ IV Conclusions
Concluding, we have introduced and experimentally validated a new explanatory framework for complex behavior observed in disordered ferroelectric crystals that support superlattices, such as KTN. Emergent phenomena are found to be a consequence of the competition between two basic topological defects that emerge from the spontaneous breaking of discrete inversion symmetry and volume-charge screening: domain walls parallel to the principal crystal axes, across which the polarization flips by 180 degrees, and slanted at 45 degrees, across which the polarization abruptly rotates by 90 degrees [89, 16, 17]. Frustration is then triggered by the interplay between the two corresponding spatial scales, each with its specific dipole-dipole interaction, a Heisenberg and a DM-like interaction. The mechanism leads to vortex and metastable phases, depending on the ratio of the two underlying interaction components. The application of the real-space renormalization group methods to detect domain patterns reveals the existence of a frustrated percolation phase transition that naturally includes depinning effects, i.e., domain wall reorientation that results in the well-known Barkhausen jumps and slim-loop hysteretic phenomena, when topologically protected states are broken. Our expansion of classical frustration mechanisms, beyond spin-glass and Griffiths effects, opens a new route to explore the intriguing interplay between structure and dynamics and how this can give rise to previously unknown phases, potentially rising to a new level the design and development of innovative miniaturized noise-resistant high-performance information and energy storage devices.
@@ Acknowledgments
L. F. and E. DelRe acknowledge support from PNRR MUR Project No. PE0000023-NQSTI, PRIN 2022 MUR Project No. 20223T577Z, Sapienza-Ricerca di Ateneo Projects. P.V. acknowledges the Spanish Ministry of Research and Innovation and Agencia Estatal de Investigación (AEI), MICIN/AEI/10.13039/501100011033, for financial support through Project PID2023-149174NB-I00, funded also by European Regional Development Funds, as well as Ref. PID2020-113681GB-I00. A. J. A. acknowledges support from the Ministry of Science, Technology and Space (IL Grant No. 4698). P.V. acknowledges the influence of the late Daniel Amit, whose vision and insights underscored the ethical responsibilities of scientific research in the face of armed conflict, systematic violations of human rights, especially regarding the ongoing genocide against the Palestinian people. We thank A. Gabrielli for extremely valuable discussions and suggestions on earlier manuscript versions.
@@ Appendix B Exploring the Metastable Phase
For completeness, we present numerical results for the temporal evolution of the polarization order parameter, $P$. Our results reveal the existence of such a metastable regime placed between the vortex and the ferroelectric phase. This is characterized, as shown in Fig. 6, by broad quasi-periodic temporal oscillations of $P$.
@@
Anomalously large sampling times would be required to extract good statistics for the actual mean values and variances. This collective behavior is a straightforward manifestation of the partial ordering of the system, being reminiscent of other types of non-ergodic behavior in structured systems [80, 44] but with a completely different origin.
@@ Appendix C Experimental Setup
Laser light from a doubled 30 mW Nd:YAG laser (wavelength $\lambda=2\pi/k_{0}=532$ nm) is made to propagate along the z-axis through the zero-cut KTN:Li crystal sandwiched in between two crossed polarizers. The sample is biased by a time-constant electric field E along the x-axis. The crystal temperature T is set by a current-controlled Peltier junction in contact with one of the y facets. Light from the crystal output facet is imaged using a moveable spherical lens L1 (of focal length 5050mm) onto a CMOS camera. The cross-polarizer technique implemented utilizes the birefringence of polar clusters to alter the polarization state of light, allowing optical observation of microscopic cluster dynamics at micrometer scales. The experimental setup is illustrated in Fig. 8.
@@ Appendix D Material
The sample is a zero-cut optical quality polished lithium-enriched solid solution of potassium-tantalate-niobate (KTN: Li) with an average composition $K_{0.964}$Li0.036Ta0.60Nb0.40O3. Its dimensions are 2.33(x) $\times$ 1.96(y) $\times$ 2.03(z) mm. The unit cell manifests random substitutions, a compositional disorder that, as a consequence of the structural flexibility typical of perovskites, leads to locally modified polarizabilities and temperature-dependent nanoscale dipolar structures (nanodisordered ferroelectricity). The result is a modified ferroelectric behavior dominated by the so-called polar nanoregions (PNRs), characterized by dielectric dispersion and out-of-equilibrium behavior (relaxor ferroelectricity) [3]. In our present case, this disorder is itself not homogeneous, manifesting a spatially periodic micrometric oscillation along a specific crystal axis. This is because the sample is grown into a bulk through the top-seeded method, a technique that entails a slight time oscillation in the temperature of the solidifying melt that, in turn, translates into an approximatively periodic 7.5 $\mu$ m pattern along the growth (or pull) axis, which causes in ferroelectrics an equally patterned $T_{c}$. This pattern conditions the nanoscale dipolar structures that, at the room-temperature Curie point $T_{c}$= 295K, form an orderly multi-domain protected regular state, the Supercrystal (SC) [8, 57, 56, 6]. SCs are thought to originate in a regular lattice of interlocked polarization vortices, a generalization of the spontaneous polarization closed-flux vortex domain pattern dominated by a triangular tiling, compatible with available experimental results and phase-field simulations [6]. This arrangement effectively screens the polarization charge and elastic stress [8, 90]. SC phase also manifests striking optical properties, such as giant broadband refraction (GR) [53, 55] and constraint-free wavelength conversion [54].